OpenStudy (unklerhaukus):

\[ x \divideontimes x =1\] \[x=?\]

OpenStudy (anonymous):

what is that sign? :S

OpenStudy (unklerhaukus):

divide on times

OpenStudy (unklerhaukus):

a bit like ±

OpenStudy (anonymous):

oh! thanks..

OpenStudy (unklerhaukus):

so there can only be two solutions right?

OpenStudy (unklerhaukus):

come on someone , take a guess,,,

OpenStudy (diyadiya):

\[x \div x=1\] =D

OpenStudy (diyadiya):

\[x \times x =x^2\] this is all i know

OpenStudy (unklerhaukus):

\[1 \divideontimes 1 =1\] True because \[1 \div 1 =1, \qquad and \qquad 1 \times 1=1\]

OpenStudy (unklerhaukus):

what is the other solution?

OpenStudy (anonymous):

I think there is only one solution.

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

No, can't be....

OpenStudy (diyadiya):

No 0/0 = indeterminate

OpenStudy (unklerhaukus):

\[\frac{0}{0} ≠ 0\]

OpenStudy (anonymous):

Yes, I never thought of that

OpenStudy (anonymous):

I agree with jasonchutko, only one solution.

OpenStudy (unklerhaukus):

what about \[-1\]

OpenStudy (diyadiya):

Oh cool!

OpenStudy (anonymous):

-1 * -1 = 1

OpenStudy (anonymous):

Very true...

OpenStudy (diyadiya):

i've never seen that sign before

OpenStudy (anonymous):

Me too..

OpenStudy (anonymous):

So the answer is one?

OpenStudy (diyadiya):

1 and -1

OpenStudy (unklerhaukus):

\[x=±1\]

OpenStudy (anonymous):

But -1 * -1 = 1!!

OpenStudy (unklerhaukus):

which satisfies the original equation \[x \divideontimes x =1\]

OpenStudy (anonymous):

Oh, yes I thought it had to be equal to the same number ... my mistake.

OpenStudy (unklerhaukus):

there is only one solution to \[x \divideontimes x =x\]\[x=1\]